# GeoLayer: Towards Low-Latency and Cost-Efficient Geo-Distributed Graph Stores with Layered Graph

### v1

@@ Abstract
Index Terms:
Geo-distributed cloud storage, graph pattern request, replica placement, request routing
@@ I Introduction
An example is the infrastructure of global payment platforms [6, 7], which supports a range of regional and global financial business scenarios: (1) Transaction graphs and account relationship networks [8] continuously generated within specific regions are maintained by local DCs to ensure low-latency processing of regional operations such as fund transfers and credit-risk assessment [9, 10]. (2) These platforms also support cross-border transactions and provide access to globally distributed financial services [11, 12], ensuring secure operations by tracking and monitoring transaction paths across regions. (3) At the global scale, these platforms conduct graph analysis to deliver personalized financial services [13, 14].
@@
Graph data access exhibits substantial pattern diversity. Queries often attach variables on vertex and edge attributes to different parts of the query pattern [31]. In practice, the attribute space is highly rich. For instance, Ant Group’s Alipay platform [32] records dozens to hundreds of key attributes for each transaction [33, 34] (e.g., counterparty information, amount, timestamp, payment method, risk score). This richness yields highly diverse query outcomes under varying constraints. Furthermore, the variability of source and destination vertices specified by different queries further amplifies the diversity of access patterns and the unpredictability of access behaviors. In addition, data accesses frequently overlap across different patterns, i.e., the same data appears in multiple patterns. Such overlap introduces significant intra-pattern variation in access frequency, further exacerbating the limitations of traditional replica placement strategies that operate at the granularity of entire patterns.
@@ II Background
(1) Online requests typically involve both read and write operations over graph patterns. For read requests, query patterns commonly involve graph pattern matching (often join-heavy on many-to-many edges with group-by, aggregation, and sorting operators) and finding paths (e.g., cheapest/shortest or weighted paths) [45]. Queries attach variables about vertex/edge attributes on different parts of the pattern [31]. For example, in a funds-transfer scenario, a read request retrieves transfers during a given period from a blocked account to an unblocked account, with variables covering the vertex type Account, vertex attribute isBlocked, edge type Transfer, and edge attribute date. For write requests, the pattern performs insert/update/delete on involved vertices, edges, or attributes. For instance, a scheduled transfer inserts a Transfer edge from the sender to the receiver and updates the corresponding account balance within the specified time window. In online mode, graph pattern requests are primarily used in interactive applications that demand fast responses, imposing strict latency requirements. To meet this, an optimized replica placement strategy can provide low-latency access.
@@
Data replication and request routing. As discussed in the online mode, a request targets a graph pattern, such as retrieving all transfers within a given time window from blocked to unblocked accounts. Because the source and destination accounts (and their related transactions) may reside in different DCs, low-latency access is achieved by on-demand replicating the pattern-relevant items to the DC closest to the user and routing the request locally whenever possible. During workload execution, access logs are continuously collected to capture per-item request frequencies, which are then used to guide adaptive replication decisions [27]. In the offline mode, existing replicas remain reusable. For joint analytics, request routing selects an optimal layout over the available replicas, favoring in-place computation and locality-preserving join, to reduce inter-DC data transfer.
TABLE III: Summary of notations.
Symbol Definition
II Set of data items
$D$ Set of geo-distributed DCs
$P$ Set of read request patterns
$s_{x}$ Size of data item $x$
$c^{store}_{d}$ Storage cost at DC $d$ per unit size
$c^{read}_{d}$ Cost for DC $d$ to process a read (GET) request
$c^{write}_{d}$ Cost for DC $d$ to process a write (PUT) request
$c^{net}_{yd}$ Network transmission cost per unit size from DC $y$ to DC $d$
$\delta_{xd}$ Binary variable indicating whether data item $x$ has a replica stored in DC $d$
$\sigma_{xyd}$ Binary variable indicating whether a read request for data item $x$ originating from DC $y$ is responded to by DC $d$
$\rho_{pyd}$ Binary variable indicating whether a read request pattern $p$ from DC $y$ is routed to DC $d$
$R_{xy}$ Read request rate for data item $x$ from DC $y$
$W_{xy}$ Write request rate for data item $x$ from DC $y$
$R_{py}$ Read request rate for pattern $p$ from DC $y$
@@ III-A System Model
We first introduce the definition of request latency, then present the $\mathsf{GeoLayer}$’s cost metrics. The major notations are summarized in Table III.
@@
$\small\text{where }\quad\Delta l^{p}_{y}=\frac{\max\limits_{\{y,d\in D\mid\rho_{pyd}=1\}}l^{p}_{yd}-\min\limits_{\{y,d\in D\mid\rho_{pyd}=1\}}l^{p}_{yd}}{\min\limits_{\{y,d\in D\mid\rho_{pyd}=1\}}l^{p}_{yd}}.$
@@ III-B Optimization Problem
s.t. $\small\text{(a)}\quad\sigma_{xyd}\leq\delta_{xd},\ \sum_{d\in D}\sigma_{xyd}=1,$
$\small\text{(b)}\quad\rho_{pyd}\leq\delta_{x^{*}d},\ \forall x^{*}\in I_{p},$
$\small\text{(c)}\quad\frac{1}{|I|}\sum_{x\in I}\sum_{y\in D}R_{xy}\sum_{d\in D}\sigma_{xyd}\cdot l^{x}_{yd}\leq\Gamma_{\max},$
$\small\text{(d)}\quad\max_{\{y,d\in D\mid\rho_{pyd}=1\}}l^{p}_{yd}\leq\eta_{p}\cdot\Gamma_{\max},$
$\small\text{(e)}\quad\delta_{xd},\ \sigma_{xyd},\ \rho_{pyd}\in\{0,1\},\ \forall x\in I,\forall y,d\in D,\forall p\in P.$
@@ IV Latency-aware layered graph
Overview. The optimization problem is a multi-objective, multi-constraint combinatorial challenge. Due to the vast decision space and strict constraints, deriving a direct global solution is computationally infeasible. To tackle this, we propose a layered graph architecture that decomposes the global optimization problem into a hierarchical tree of subproblems based on latency-aware layering.
Specifically, patterns involving a large number of cross-partition edges inherently trigger inter-DC data accesses [53], with request latencies varying significantly due to network heterogeneity. We organize these cross-partition edges into multiple layers based on ascending request latency. The inter-layer latency differences exhibit a monotonic progression, enabling the effective alignment of latency-sensitive patterns with appropriate layers. Within each layer, a set of edges connects lower-layer graph regions with similar response latencies, forming latency-homogeneous clusters. The connections among these layer-wise edge sets and clusters form a hierarchical tree. By confining the optimization process to the branches of this tree, we effectively reduce decision complexity while alleviating the inherent coupling between decision variables in joint optimization.
Layered Graph Construction. Given a connected graph and its geo-distributed partitioning, the layered graph is constructed by combining three components: (1) the base layer, denoted as $Layer_{0}$, which consists of per-DC subgraphs $\{G_{d}=(V_{d},E_{d})\}_{d\in D}$; (2) upper-layer bridge graphs ($Layer_{i},1\leq i\leq h$), denoted as $Layer_{i}:G^{B}_{i}=(V^{B}_{i},E^{B}_{i})$ which capture inter-partition connectivity at different latency levels; and (3) the inter-layer links that connect bridge graphs across adjacent layers.
@@ V-A Design Overview
Building on the layered graph, our replica placement scheme localizes decision-making within clusters at each layer. The process begins with a sinking step that assigns each pattern to the latency-compliant layer. To address the variation in access frequencies across overlapping regions of patterns, we first apply a gain function to determine whether a pattern should undergo full intra-cluster replication or decomposition. For patterns deemed decomposable, we split them along their overlapping regions and let these region‐segments competitively bid for placement across the cluster’s bridge subgraphs. To quantify the affinity between each bridge subgraph and these segments, we introduce a directed heat diffusion model, which guides the placement of pattern segments within each cluster. This process cascades recursively through the layers until replicas are placed in individual DCs.
1
Output: Final pattern distribution $\mathbb{P}$
2
3 Function $\textnormal{{ReplicaSink}}()$
4
5 for $k\leftarrow h$ to 11 do
6     while $\mathbb{P}^{(k)}\neq\emptyset$ do
7        $p\leftarrow\mathbb{P}^{(k)}.\text{pop}()$
8        if $\eta_{p}\cdot\Gamma_{\max}\in(t_{2k-1},t_{2k})$ then
9           $\textnormal{{PlaceReplica}}(p,\mathbb{P}^{(k)},L_{k})$
10
11        else
12          //  Latency violation - sinking
13           $\textnormal{{FullReplicate}}(p,L_{k-1})$
14           $\textnormal{{RecordPlacement}}(p,L_{k-1})$
15           $\mathbb{P}^{(k-1)}\leftarrow\mathbb{P}^{(k-1)}\cup\{p\}$
16
17
18
19 return $\mathbb{P}$
Algorithm 1 Pattern Replica Sinking Process
@@ V-B Replica Placement Built on Layered Graph
Sinking of Pattern Replicas. Graph pattern requests impose varying latency requirements, as reflected in constraints (d) of Equation (6). Because the layered graph is organized by increasing request latency, whenever a pattern’s latency bound cannot be met at layer $L_{k}$, a complete replica of that pattern must be sunk to every bridge subgraph in the corresponding clusters of $L_{k-1}$. In other words, the components of the pattern cannot be split across different bridge subgraph regions in layer $L_{k-1}$, as such dispersed placement would fail to meet the latency requirements for remote access between components within the pattern. This sinking process continues until the pattern reaches a layer that satisfies its latency requirements. Once a layer does meet the pattern’s latency constraint, $\mathsf{GeoLayer}$ allows the pattern’s data items to be decomposed and distributed across the appropriate regions in all lower layers, enabling finer-grained cost optimization. At that point, the pattern is placed at this layer in preparation for subsequent replica placement optimization. Algorithm 1 describes the entire process. The input pattern set $P$ is derived from historical access logs and carries request frequency and latency requirement features. The layer set $L$ is constructed using a fixed-threshold bucketing scheme (e.g., 100 ms per interval). The mechanism of sinking replicas enforces strict latency guarantees while enabling a trade-off between the cost of remote access and the benefits of local caching.
@@
Intuitively, in real-world graphs, core regions (e.g., celebrities in social networks) sustain high access frequencies, and their influence propagates along edges to neighboring regions, elevating the neighbors’ access rates. Moreover, overlapping influences from multiple cores superimpose at common neighbors, making their accesses even more frequent. This observation suggests viewing access as heat: vertices as heat sources, edges conduct heat, and the heat field cools unless reheated by new requests.

### v3

@@ I Introduction
An example is the infrastructure of global payment platforms [7, 8], which supports a range of regional and global financial business scenarios: (1) Transaction graphs and account relationship networks [9] continuously generated within specific regions are maintained by local DCs to ensure low-latency processing of regional operations such as fund transfers and credit-risk assessment [10, 11]. (2) These platforms also support cross-border transactions and provide access to globally distributed financial services [12, 13], ensuring secure operations by tracking and monitoring transaction paths across regions. (3) At the global scale, these platforms conduct graph analysis to deliver personalized financial services [14].
@@
Graph data access exhibits substantial pattern diversity. Queries often attach variables on vertex and edge attributes to different parts of the query pattern [31]. In real-world financial systems, such as Ant Group’s Alipay platform [32], each transaction records dozens of attributes that are directly relevant to queries [33, 34] (e.g., counterparty information, amount, timestamp, payment method, risk score). These attributes collectively define rich graph patterns that closely reflect real query workloads. Furthermore, the variability of source and destination vertices specified by different queries further amplifies the diversity of access patterns and the unpredictability of access behaviors. In addition, data accesses frequently overlap across different patterns, i.e., the same data appears in multiple patterns. Such overlap introduces significant intra-pattern variation in access frequency, further exacerbating the limitations of traditional replica placement strategies that operate at the granularity of entire patterns.
@@ II Background
(1) Online requests typically involve both read and write operations over graph patterns. For read requests, query patterns commonly involve graph pattern matching (often join-heavy on many-to-many edges with group-by, aggregation, and sorting operators) and finding paths (e.g., cheapest/shortest or weighted paths) [45, 46]. Queries attach variables about vertex/edge attributes on different parts of the pattern [31]. For example, in a funds-transfer scenario, a read request retrieves transfers during a given period from a blocked account to an unblocked account, with variables covering the vertex type Account, vertex attribute isBlocked, edge type Transfer, and edge attribute date. For write requests, the pattern performs insert/update/delete on involved vertices, edges, or attributes. For instance, a scheduled transfer inserts a Transfer edge from the sender to the receiver and updates the corresponding account balance within the specified time window. In online mode, graph pattern requests are primarily used in interactive applications that demand fast responses, imposing strict latency requirements. To meet this, an optimized replica placement strategy can provide low-latency access.
@@
Data replication and request routing. As discussed in the online mode, requests target graph patterns, such as retrieving all transfers within a given time window from blocked to unblocked accounts. Because the source and destination accounts (and their related transactions) may reside in different DCs, low-latency access is achieved by on-demand replicating the pattern-related items to the DC closest to the user and routing the request locally whenever possible. During workload execution, access logs are continuously collected to capture per-item request frequencies, which are then used to guide adaptive replication decisions [27]. In the offline mode, existing replicas remain reusable. For joint analytics, request routing selects an optimal layout over the available replicas, favoring in-place computation and locality-preserving join, to reduce inter-DC data transfer.
@@ III-A System Model
We first introduce the definition of request latency, then present the $\mathsf{GeoLayer}$’s cost metrics.
@@
$\small\text{where }\qquad\Delta l^{p}_{y}=\frac{\max\limits_{\{y,d\in D\mid\rho_{pyd}=1\}}l^{p}_{yd}-\min\limits_{\{y,d\in D\mid\rho_{pyd}=1\}}l^{p}_{yd}}{\min\limits_{\{y,d\in D\mid\rho_{pyd}=1\}}l^{p}_{yd}}.$
@@ III-B Optimization Problem
s.t. $\small\text{(a)}\qquad\sigma_{xyd}\leq\delta_{xd},\ \sum_{d\in D}\sigma_{xyd}=1,$
$\small\text{(b)}\qquad\rho_{pyd}\leq\delta_{x^{*}d},\ \forall x^{*}\in I_{p},$
$\small\text{(c)}\qquad\frac{1}{|I|}\sum_{x\in I}\sum_{y\in D}R_{xy}\sum_{d\in D}\sigma_{xyd}\cdot l^{x}_{yd}\leq\Gamma_{\max},$
$\small\text{(d)}\qquad\max_{\{y,d\in D\mid\rho_{pyd}=1\}}l^{p}_{yd}\leq\eta_{p}\cdot\Gamma_{\max},$
$\small\text{(e)}\qquad\delta_{xd},\ \sigma_{xyd},\ \rho_{pyd}\in\{0,1\},\ \forall x\in I,\forall y,d\in D,\forall p\in P.$
@@ IV Latency-aware layered graph
Overview. The optimization problem is a multi-objective, multi-constraint combinatorial task for which direct global optimization is computationally intractable. We address this by constructing a latency-aware layered graph that decomposes the global problem into a hierarchical tree of subproblems. Patterns involving a large number of cross-partition edges inherently trigger inter-DC data accesses [54], with request latencies varying significantly due to network heterogeneity. We organize these cross-partition edges into multiple layers based on ascending request latency. By construction, inter-layer latencies progress monotonically. Within each layer, a set of edges with similar response latencies groups the connected lower-layer graph regions into latency-homogeneous clusters. The connections among these layer-wise edge sets and clusters form a hierarchical tree. Restricting placement and routing decisions to the branches of this tree substantially reduces the decision space, mitigates variable coupling, and provides a unified, stable structure for subsequent overlap-centric replica placement (§V) and stepwise layered routing (§VI).
Layered Graph Construction. Given a connected graph $G=(V,E)$ and its geo-distributed partitioning, the layered graph is constructed by combining three components: (1) the base layer, denoted as $Layer_{0}$, which consists of per-DC subgraphs $\{G_{d}=(V_{d},E_{d})\}_{d\in D}$; (2) upper-layer bridge graphs ($Layer_{i},1\leq i\leq h$), denoted as $Layer_{i}:G^{B}_{i}=(V^{B}_{i},E^{B}_{i})$ which capture inter-partition connectivity at different latency levels; and (3) the inter-layer links that connect bridge graphs across adjacent layers.
@@ V-A Design Overview
Building on the layered graph, our replica placement scheme localizes decision-making within clusters at each layer, operating at the pattern/overlap-region granularity. The process begins with a sinking step that assigns each pattern to the latency-compliant layer. To address the variation in access frequencies across overlapping regions of patterns, we first apply a gain function to determine whether a pattern should undergo full intra-cluster replication or decomposition. For decomposable patterns, we split them along their overlapping regions and let these region‐segments competitively bid for placement across the cluster’s bridge subgraphs. To quantify the affinity between each bridge subgraph and these segments, we introduce a directed heat diffusion model, which guides the placement of pattern segments within each cluster. This process cascades recursively through the layers until replicas are placed in individual DCs. By decomposing patterns at the overlap-region granularity, we weaken co-location coupling among pattern-related data and restrict placement decisions to latency-homogeneous clusters. Consequently, routing cost depends only on whether a cluster is covered, rendering the objective cluster-separable and enabling a place-first, then best-response routing (§VI). Due to space constraints, we defer the full implementation details to [1].
@@ V-B Replica Placement Built on Layered Graph
Sinking of Pattern Replicas. Graph pattern requests impose varying latency requirements, as reflected in constraints (d) of Equation (6). Because the layered graph is organized by increasing request latency, whenever a pattern’s latency bound cannot be met at layer $L_{k}$, a complete replica of that pattern must be sunk to every bridge subgraph in the corresponding clusters of $L_{k-1}$. In other words, the components of the pattern cannot be split across different bridge subgraph regions in layer $L_{k-1}$, as such dispersed placement would fail to meet the latency requirements for remote access between components within the pattern. This sinking process continues until the pattern reaches a layer that satisfies its latency requirements. Once a layer does meet the pattern’s latency constraint, $\mathsf{GeoLayer}$ allows the pattern’s data items to be decomposed and distributed across the appropriate regions in all lower layers, enabling finer-grained cost optimization. At that point, the pattern is placed at this layer in preparation for subsequent replica placement optimization. The mechanism of sinking replicas enforces strict latency guarantees while enabling a trade-off between the cost of remote access and the benefits of local caching.
@@
Intuitively, in real-world graphs, core regions (e.g., celebrities in social networks) attract high request rates, and their influence propagates along edges to neighboring regions, elevating neighbors’ access frequencies. Overlapping influence from multiple cores superimposes at common neighbors, making their accesses even more frequent. This motivates a heat view of access: vertices act as heat sources, edges conduct heat, and the heat field dissipates unless reheated by new requests.
@@
The convergence of the DHD model is formally established, and the detailed proof is deferred to  [1] due to the page limit.
$\text{where}\qquad\small\Delta\hat{C}^{(A)}_{xy}=R_{xy}\cdot\left(\lambda_{1}\sum_{j=1,BS_{j}\in\mathcal{C}}^{n}\sigma_{xyj}\right).$
(a) Pattern overlap
(b) Overlap-centric diffusion
To proactively guard against sudden surges in access frequency, we use the DHD model to identify and pre-cache these high-heat regions. First, the DC issues a global unobserved edge request to upper layers to acquire complete topological connectivity of externally cached regions. Concurrently, it initializes a high-heat vertex set $\mathcal{O}=\{v\in V\mid\mathcal{H}_{v}^{0}\geq\theta\}$, treating these vertices as external heat sources as defined in Equation (9). The system then iteratively updates heat states using Equation (10) until it reaches steady state (see Equation (12)). At equilibrium, we extract the high-heat subgraph $G^{H}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\{v\in V\mid\mathcal{H}_{v}^{*}\geq\theta\}$ and $\mathcal{E}=\{(u,v)\mid u,v\in\mathcal{V}\}$. This subgraph is subsequently pre-cached in the DC. The guarantee of the lower bound on pre-caching probability is deferred to [1] due to space limitations.
Update Maintenance. We adopt a hybrid maintenance strategy that combines incremental updates with periodic refreshes to handle dynamic data changes. The layered graph is a persistent, offline-built structure, with replicas cached within DCs to anticipate future accesses. During periodic maintenance, newly inserted items that have been accessed are captured by access logs, thereby materializing request patterns. These patterns are then fed into the overlap-centric replica placement algorithm, which reacts to these pattern changes and updates replica locations accordingly.
When a vertex or edge is deleted within a DC, maintenance starts at ${Layer}_{0}$. We trigger a bottom-up cleanup: the DC handling the delete operation updates its local state, and the removal propagates upward along bridge subgraphs to affected upper-layer clusters until all replicas are removed. In addition, the online replica eviction mechanism gradually evicts replicas as deleted data cools down, assisting the deletion process.
VI Stepwise Layered Routing
Figure 5: Request routing under the online query mode.
Figure 6: Request routing under the offline query mode.
VII Experimental evaluation
VII-A Experimental Setup
Figure 7: Request latency.
Figure 8: Comparison of cost metrics.
Figure 9: Optimality gap.
Figure 10: Execution time.
Figure 11: Sensitivity to network heterogeneity.
Figure 12: Impact of parameter $\theta$.
Datasets. Our evaluation employs three distinct datasets: a synthetic social network generated using the LDBC-SNB benchmark with a scale factor (SF) of 10, and two real-world graph datasets, UK-2005 [58] and Twitter-2010[59]. Table III provides a detailed summary of the statistics for each dataset.
TABLE III: Dataset Description
Graph Vertices Edges Abbreviation
LDBC-SNB-SF10 [60] 36,423,181 231,371,359 SNB
UK-2005 [58] 39,459,925 936,364,282 UK
Twitter-2010 [59] 41,652,230 1,468,364,884 TW
[…]
Figure 13: Runtime of graph analysis.
Figure 14: WAN cost of graph analysis.
Figure 15: Migration ratio.
VII-B Effectiveness Analysis in Online Mode
[…]
We also report the algorithm execution time, as shown in Figure 12. Random-3 and Top-3 exhibit lower computational complexity due to their simple heuristic strategies. ADP incurs higher overhead as it relies on hypergraph partitioning and multiple rounds of iteration. Although $\mathsf{GeoLayer}$ involves additional complexity due to its multi-layered design, it remains efficient by leveraging cluster-level parallelism, producing results within a reasonable execution time.
[…]
Optimality Gap. To quantify the closeness of $\mathsf{GeoLayer}$ to the optimal solution, we employ an MILP solver (PuLP/CBC) [70] on the small-scale WIKI-vote dataset [71] (7,115 vertices and 103,689 edges). The setup involves four DCs, with 1,000 historical query patterns and 100 test patterns. We solve Equation (6) to obtain the global optimum ($C^{*}$) for comparison. Figure 9 reports the costs of $\mathsf{GeoLayer}$ and the global optimum across different metrics, with all values normalized to $\mathsf{GeoLayer}$ ’s total cost ($C$). We compute the overall optimality gap as: $\mathrm{Gap}=\frac{C-C^{*}}{C^{*}}\times 100\%=7.8\%$. This result shows that $\mathsf{GeoLayer}$ achieves small optimality gaps under the same constraints, indicating that our replication and routing schemes are near-optimal on small graphs, while scaling effectively to large deployments where exact optimization is intractable.
[…]
VII-C Performance Comparison in Offline Mode
[…]
WAN Cost. We evaluate WAN usage during algorithm execution by measuring the actual volume of inter-DC data transfer. As shown in Figure 15, $\mathsf{GeoLayer}$ incurs the lowest WAN cost across all test conditions. Specifically, $\mathsf{GeoLayer}$ reduces WAN cost by an average of 42.1% (up to 72.0%) compared to RAGraph, 28.1% (up to 46.8%) compared to RAGraph+, and 34.7% (up to 68.2%) compared to GrapH. The reduction in WAN usage stems from $\mathsf{GeoLayer}$’s ability to select fewer execution sites while minimizing the introduced cross–DC communication. $\mathsf{GeoLayer}$ constructs $2$ execution sites for the $\mathsf{SNB}$ graph, and $3$ sites for both the $\mathsf{UK}$ and $\mathsf{TW}$ graphs.
Migration Ratio. We also report the ratio of migrated edges under each strategy, as shown in Figure 15. $\mathsf{GeoLayer}$ achieves a migration ratio ranging from 34% to 42% across the three datasets. On the $\mathsf{SNB}$ graph, $\mathsf{GeoLayer}$ exhibits a relatively higher migration ratio, primarily due to the consolidation of data into only two execution sites to enable more efficient processing. All evaluated strategies consider network heterogeneity in geo-distributed environments and aim to improve graph analysis performance by reducing communication overhead through limited data movement.
VII-D Ablation Study
In this subsection, we quantify the contributions of the replica placement and request routing components under both online and offline workloads across three datasets. We adopt $\mathsf{RP+RR}$ (3-replica random placement + random routing) as the baseline. To isolate component benefits, we replace one or both modules: (i) $\mathsf{RP+SR}$: random placement with stepwise layered routing (routing contribution); (ii) $\mathsf{LP+RR}$: overlap-centric replica placement with random routing (placement contribution); and (iii) $\mathsf{LP+SR}$: both overlap-centric placement and stepwise layered routing (full configuration). For offline analytics, We report $\mathsf{PageRank}$ as a representative algorithm.
Figure 16: Performance gain.
[…]
VIII Related Works
Data Replication and Request Routing. A large body of work [18, 19, 20, 21, 22, 23, 24, 72, 73, 74, 75, 76, 77, 78] focuses on optimizing replica placement and request routing decisions in geo-distributed environments. These efforts address the problem from multiple perspectives, considering factors such as replica placement [22, 23, 72, 24], caching time [18], data migration [68, 69], QoS/QoE/SLA guarantees [19, 20, 74], consistency models [21, 75], security policies [76, 77], and load balancing [78, 73]. Several studies further incorporate data association into decision-making processes [25, 17, 26, 27, 28, 29, 79]. Volley [79] proposes an iterative optimization algorithm based on weighted spherical means to handle inter-data dependencies. Yu et al. [28] model data association as a hypergraph partitioning problem to guide co-location decisions. Liu et al. [27] develop a replica placement strategy based on overlapping community detection to handle associated data. These approaches commonly abstract associated data items as edges and connect them with storage nodes to form a graph structure, thereby enabling graph-based modeling of placement decisions. However, these associations are driven by job-level relationships, whereas $\mathsf{GeoLayer}$ focuses on the intrinsic structural properties of the graph.
Geo-Distributed Graph Data Processing. Many existing efforts [80, 81, 48, 82, 83, 84, 85, 47] aim to optimize the performance of graph-structured data processing in geo-distributed settings. GeoGraph [80] leverages clustered DCs and a metagraph to support graph query processing. PGPregel [48] adopts sampling and combiner techniques to enable privacy-preserving graph processing. Hosseinalipour et al. [82] propose an energy-efficient graph job assignment method for geo-distributed cloud networks. RAGraph [47] exploits graph monotonicity to design heterogeneity-aware geo-distributed graph computation strategies. These works focus on optimizing the performance of upstream computing systems, and are thus orthogonal to $\mathsf{GeoLayer}$’s design.
Layered Graph Structures. Layered designs are widely used in graph computation and analysis. CompressGraph [86] builds a two-layer traversal model via graph compression. Layph [64] partitions a graph into an upper-layer skeleton and lower-layer disjoint subgraphs to constrain the scope of incremental computation in dynamic graph processing. In road networks, hierarchical structures with shortcut edges are standard for speeding up path queries [87, 88]. These layered approaches are tailored to the computational characteristics of graph data, but do not consider heterogeneous latency on edges. In contrast, $\mathsf{GeoLayer}$ adopts a layered structure specifically designed for the latency characteristics of cross-partition edges, marking a fundamental distinction from prior work.
IX Conclusion
Appendix
We first present the theoretical analysis of $\mathsf{GeoLayer}$, followed by its concrete implementation. The major notations are summarized in Table IV.
TABLE IV: Summary of notations.
Symbol Definition
II, $D$, $P$, Sets of data items ($I$), geo-distributed DCs ($D$), and read request patterns ($P$)
$s_{x}$ Size of data item $x$
$c^{store}_{d}$, $c^{read}_{d}$, $c^{write}_{d}$, $c^{net}_{yd}$ Costs: storage cost at DC $d$ per unit size; read (GET) processing cost at DC $d$; write (PUT) processing cost at DC $d$; network transmission cost per unit size from DC $y$ to DC $d$
$\delta_{xd}$, $\sigma_{xyd}$, $\rho_{pyd}$ Binary indicators: $\delta_{xd}$ indicates whether item $x$ has a replica in DC $d$; $\sigma_{xyd}$ indicates whether a read request for $x$ from DC $y$ is served by DC $d$; $\rho_{pyd}$ indicates whether a pattern $p$ read from DC $y$ is routed to DC $d$
$R_{xy}$, $W_{xy}$, $R_{py}$ Request rates: $R_{xy}$ is the read rate for item $x$ from DC $y$; $W_{xy}$ is the write rate for item $x$ from DC $y$; $R_{py}$ is the read rate for pattern $p$ from DC $y$
$V_{d}$, $E_{d}$, $V^{B}$, $E^{B}$ Vertex set and edge set in DC $d$ ($V_{d}$, $E_{d}$); boundary-vertex set and cross-partition edge set ($V^{B}$, $E^{B}$)
$\alpha$, $\gamma$, $\beta$ Diffusion intensity; heat decay rate; external heat-source injection coefficient.
BSBS, $\mathcal{C}$ Bridge subgraph and a cluster
IX-A Convergence of DHD Model